# Robust Design and Evaluation of

### v5

@@ 3 Robust Design of Predictive Algorithms
In this section, we propose simple nonparametric regression estimators for the bounds on the conditional likelihood $\mu^{*}(\cdot)$ under Equation (6). Building on recent work studying the estimation of heterogeneous treatment effects in causal inference (e.g., KunzelEtAl(19); NieWager(20); Kennedy(22)-towardsoptimal), our estimators are based on cross-fitting and incorporate doubly-robust style bias correction using efficient influence functions. Since the bounds on the conditional likelihood are an infinite-dimensional parameter, we establish finite-sample bounds on the integrated mean square error (MSE) convergence of our estimators using a novel oracle inequality for pseudo-outcome regression procedures.
To simplify exposition in the main text, we develop our nonparametric regression estimator assuming the researcher specifies nonparametric outcome bounds (Section 2.3.1) and illustrate it with two folds. The analysis for multiple folds and cross-fitting is straightforward. We extend our estimators to alternative choices of bounding functions, such as proxy outcome bounds in Appendix C.3 and instrumental variable bounds in Appendix C.4.
To estimate the bounds on the conditional likelihood, we use the first fold to construct estimates of the nuisance functions $\hat{\eta}=\left(\hat{\mu}_{1}(\cdot),\hat{\pi}_{1}(\cdot)\right)$. On the second fold, we construct the pseudo-outcomes $\phi_{\mu,i}(\hat{\eta})+\left(\overline{\Gamma}-1\right)\phi_{\pi\mu,i}(\hat{\eta})$ for the upper bound and $\phi_{\mu,i}(\hat{\eta})+\left(\underline{\Gamma}-1\right)\phi_{\pi\mu,i}(\hat{\eta})$ for the lower bound, where $\phi_{\mu,i}(\eta):=\mu_{1}(X_{i})+\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))$ is the uncentered efficient influence function for $\mathbb{E}\left[\mu_{1}(X_{i})\right]$ and $\phi_{\pi\mu,i}(\eta):=\left((1-D_{i})-\pi_{0}(X_{i})\right)\mu_{1}(X_{i})+\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))\pi_{0}(X_{i})+\pi_{0}(X_{i})\mu_{1}(X_{i})$ is the uncentered efficient influence function for $\mathbb{E}\left[\pi_{0}(X_{i})\mu_{1}(X_{i})\right]$. These can be derived using standard arguments (e.g., Hines_2022). We finally regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure in the second fold. This yields our estimators $\widehat{\underline{\mu}}(\cdot)$, $\widehat{\overline{\mu}}(\cdot)$.
@@ 3.1 Bound on integrated mean square error convergence
We analyze the integrated MSE convergence rate of our proposed estimators, comparing them against an infeasible oracle nonparametric regression with access to the true nuisance functions $\eta$. The oracle estimates the conditional probability bounds by regressing the true pseudo-outcomes $\phi_{\mu,i}(\eta)+\left(\overline{\Gamma}-1\right)\phi_{\pi\mu,i}(\eta)$ for the upper bound and $\phi_{\mu,i}(\eta)+\left(\underline{\Gamma}-1\right)\phi_{\pi\mu,i}(\eta)$ for the lower bound on $X_{i}$ in the second fold using the same nonparametric regression procedure as specified by the researcher. Denote these oracle estimators by $\widehat{\underline{\mu}}_{oracle}(\cdot),\widehat{\overline{\mu}}_{oracle}(\cdot)$.
Under a mild stability condition on the second-stage nonparametric regression procedure chosen by the researcher, the integrated MSE of our feasible estimator is bounded by that of the oracle estimator plus a smoothed, doubly-robust remainder.
@@
Let $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ denote the second-stage pseudo-outcome regression estimator. Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), and $\mathbb{P}(\epsilon\leq\hat{\pi}_{1}(X_{i})\leq 1-\epsilon)=1$ for some $\epsilon>0$. Define $\tilde{R}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, and $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$. Then,
@@
Proposition 3.1 establishes that the integrated MSE of our feasible estimators of the conditional probability bounds can be no larger than that of an infeasible oracle nonparametric regression that has access to the true nuisance functions plus the $L_{2}(\mathbb{P})$-norm of a smoothed remainder term $\tilde{R}(x)$ that depends on a doubly-robust-style product of nuisance function estimation errors. The first-step estimation of the nuisance functions therefore only affects the integrated MSE of our estimator through this remainder term.
To prove this result, we establish a novel oracle inequality on the $L_{2}(\mathbb{P})$-error of pseudo-outcome regression procedures (Lemma B.1), extending Kennedy(22)-towardsoptimal’s pointwise analysis of pseudo-outcome regression procedures. This result may be of independent interest. The $L_{2}(\mathbb{P})$-stability condition (Assumption B.1) on $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ required for the oracle inequality is quite mild in practice. It is satisfied by a variety of generic linear smoothers such as linear regression, series regression, nearest neighbor matching, random forest models, and several others (Proposition B.1). Proposition 3.1 is therefore agnostic to both the researcher’s choice of ${\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]}$ and nuisance function estimators in this sense. We conjecture that there are estimators that are not linear smoothers that satisfy the $L_{2}(\mathbb{P})$-stability condition, but leave this to future work.
Furthermore, Proposition 3.1 can be applied in settings where the nuisance functions and bounding functions satisfy additional smoothness or sparsity conditions, and for particular choices of the second-stage regression estimator and nuisance function estimators. Known results on the mean-squared error convergence rates of nonparametric regression procedures can then be used to analyze $\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|$ and $\|\tilde{R}(\cdot)\|$, and give specific rates of convergence in terms of the primitives of the setting like the underlying smoothness or sparsity assumptions and the sample size following in the spirit of recent work on the optimal estimation of heterogeneous treatment effects (Kennedy(22)-towardsoptimal; KennedyEtAl(23)-Minimax).
@@ 3.2 Regret bound for plug-in decision rules
Researchers often apply a simple threshold rule to an estimated predictive algorithm in order to make decisions, as mentioned earlier. For example, KLLLM(18); Rambachan(21); AngelovaEtAl(22) compare outcomes under alternative decisions rules that threshold a predictive algorithm in pretrial release. We next show that plug-in decision rules based on our nonparametric estimators of the conditional probability bounds produce robust decisions.
Since the expected welfare of any decision rule $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ is only partially identified (Lemma 2.2), we evaluate decision rules by comparing their worst-case expected welfare (e.g., Manski(07)). [^12] Christensen2023optimal adopt an analogous criterion in analyzing optimal decision rules in parametric or semiparametric statistical models in which some payoff relevant parameter is partially identified. The optimal max-min decision rule is $d^{*}(\cdot)\in\arg\max_{d(\cdot)\colon\mathcal{X}\rightarrow[0,1]}\underline{U}(d)$, and notice that we place no restrictions on the class of decision rules.131313 This contrasts with recent work on statistical treatment assignment rules such as KitagawaTetenov(18); AtheyWager(21); KallusZhou(21), which typically compares estimated decisions rules against the best decision rule in some restricted policy class. An immediate consequence of Lemma 2.2 is that the optimal, max-min decision rule is a simple threshold rule on a particular welfare-weighted average of the conditional probability bounds
$d^{*}(X_{i})=1\{\widetilde{\mu}^{*}(X_{i})\leq u_{1,0,i}+u_{0,0,i}\}$
for $\widetilde{\mu}^{*}(x)=(u_{1,1,i}+u_{1,0,i})\overline{\mu}^{*}(x)+(u_{0,0,i}+u_{0,1,i})\underline{\mu}^{*}(x)$. Of course, the true conditional probability bounds are not known, and so we instead consider a feasible decision rule that simply plugs in our nonparametric estimators into the optimal threshold rule. We denote this plug-in decision rule as $\hat{d}(X_{i})$.
The regret of the plug-in decision rule is $R(\hat{d})=\underline{U}(d^{*})-\underline{U}(\hat{d})$, which compares the worst-case expected welfare of the max-min optimal decision rule against what is attained by the plug-in decision rule. Our next result bounds the regret of the plug-in decision rule.
@@
$R(\hat{d})^{2}\leq 2\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|+2\|\widehat{\underline{\mu}}_{oracle}(\cdot)-\underline{\mu}^{*}(\cdot)\|+4\epsilon^{-1}\sqrt{2}\|\tilde{R}(x)\|+o_{\mathbb{P}}(R_{oracle}).$
The squared regret of the plug-in decision rule is bounded above by the integrated mean square error of the oracle procedure for the conditional probability bounds plus again a smoothed, doubly-robust remainder term that depends on the product of nuisance function errors. Proving this result involves establishing that the regret of the plug-in decision rule is bounded above by the integrated mean square error of our estimators for the true conditional probability bounds. We can then apply Proposition 3.1 to show that the worst-case regret of the plug-in decision rule therefore converges to zero whenever the integrated MSE of the oracle converges to zero (and so does the integrated MSE of the our estimators). Plug-in decision rules based on our nonparametric estimators lead to robust decisions in this formal sense.
@@ 4 Robust Evaluation of Predictive Algorithms
In this section, we construct estimators for the bounds on the performance of a predictive algorithm $s(\cdot)$ under Equation (6). Our estimators are consistent for the bounds on the identified set as the sample size grows large, and we characterize their convergence rates in terms of the errors in the first-step estimation of nuisance parameters.
To simplify exposition in the main text, we again assume the researcher specifies nonparametric outcome bounds. We extend to alternative choices of bounding functions, such as proxy outcome bounds and instrumental variable bounds in Online Appendices C.3-C.4 respectively.
@@ 4.1 Bounds on positive class performance
We first estimate the bounds on positive class performance. As shown in Lemma 2.1, the bounds on positive class performance are given by the optimal values of linear-fractional programs, and consequently they are not pathwise differentiable without further assumptions. Standard arguments for constructing debiased estimators based on the von Mises expansion therefore do not directly apply (e.g., see Section 5 of Kennedy(22)-IFreview). However, we show that the structure of linear-fractional programming can nonetheless be exploited to extend the construction of debiased estimators based on our pseudo-outcome regression procedure.
We illustrate our estimator for the bounds on positive class performance using sample-splitting across three folds. Extending to multiple folds and cross-fitting is again straightforward. To estimate the bounds on positive class performance, we randomly split the data into three folds (we assume $n$ is divisible by $3$ and each fold contains $n/3$ observations for simplicity). We construct nuisance function estimates $\hat{\eta}$ using all observations in the first fold. On the second fold, we construct some nonparametric regression estimator $\hat{\delta}(\cdot)$ of the bounding functions, and we will discuss the specific choice of estimator next. Finally, on the third fold, we solve the sample analogue
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\widetilde{\delta}\in\widehat{\Delta}_{n}}\frac{\mathbb{E}_{n}\left[\beta_{0,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\beta_{0,i}\widetilde{\delta}_{i}\right]}{\mathbb{E}_{n}\left[\phi_{\mu,i}(\hat{\eta}_{-k})+\widetilde{\delta}_{i}\right]},$ (9)
where $\widehat{\Delta}_{n}:=\left\{\widetilde{\delta}\colon(\underline{\Gamma}-1)(1-D_{i})\widehat{\delta}(X_{i})\leq\widetilde{\delta}_{i}\leq(\overline{\Gamma}-1)(1-D_{i})\widehat{\delta}(X_{i})\mbox{ for }i=1,\ldots,n/3\right\}$ and $\mathbb{E}_{n}[\cdot]$ is the sample average over all observations in the third fold. In other words, we directly solve the sample analogue of linear-fractional program characterizing the bounds by plugging in estimated constraints. Since Equation (9) is a linear-fractional program, it can be equivalently expressed as a linear program by applying the Charnes-Cooper transformation (CharnesCooper(62)). Our estimator for the bounds on positive class performance therefore involves solving a linear program (see Online Appendix C.1 for details).
Our next result characterizes the convergence rate of our estimators to the true bounds on positive class performance.

### v6

@@ 3 Robust Design of Predictive Algorithms
In this section, we propose estimators for the bounds on the conditional likelihood $\mu^{*}(\cdot)$ under Assumption 2.1. We build on recent work studying the estimation of heterogeneous treatment effects in causal inference (e.g., KunzelEtAl(19); NieWager(20); Kennedy(22)-towardsoptimal), and our estimators use cross-fitting and doubly-robust bias corrections. Since the bounds on the conditional likelihood are an infinite-dimensional parameter, we establish finite-sample bounds on our estimator’s integrated mean square error (MSE) convergence using a novel oracle inequality for pseudo-outcome regression procedures. These convergence guarantees enable us to construct plug-in decision rules with provable regret bounds.
To simplify exposition, we develop our estimator assuming the researcher specifies observed outcome bounds (Section 2.3.1) and illustrate it with two folds. The analysis for multiple folds and cross-fitting is straightforward. We extend to proxy outcome bounds in Online Appendix D.1 and instrumental variable bounds in Online Appendix D.2.
Estimation Procedure for Observed Outcome Bounds:
Under Assumption 2.1, the unobserved conditional probability among unselected units can be decomposed as $P(Y^{*}_{i}=1|D_{i}=0,X_{i}=x)=\mu_{1}(x)+\delta(x)$, where $\mu_{1}(x)$ is the observed rate among selected units and $\delta(x)$ measures confounding. Our estimation strategy constructs pseudo-outcomes based on influence functions that combine observed outcomes among selected units with extrapolation to unselected units. The pseudo-outcomes correct for both observable selection (via inverse propensity weighting) and unobservable confounding (via the sensitivity parameters $\underline{\Gamma},\overline{\Gamma}$).
[…]
@@ 3.1 Bound on Integrated Mean Square Error Convergence
We analyze the integrated MSE convergence rate of our proposed estimators by comparing them against an infeasible oracle nonparametric regression. This oracle has access to the true nuisance functions $\eta$ and estimates the conditional likelihood bounds by regressing the true pseudo-outcomes $\phi_{\mu,i}(\eta)+\left(\overline{\Gamma}-1\right)\phi_{\pi\mu,i}(\eta)$ for the upper bound and $\phi_{\mu,i}(\eta)+\left(\underline{\Gamma}-1\right)\phi_{\pi\mu,i}(\eta)$ for the lower bound on $X_{i}$ in the second fold using the same nonparametric regression procedure specified by the researcher. We denote these oracle estimators by $\widehat{\underline{\mu}}_{oracle}(\cdot),\widehat{\overline{\mu}}_{oracle}(\cdot)$.
This oracle comparison is useful because it decomposes the integrated MSE of our proposed estimators into two sources. The first is approximation error from using a finite-sample nonparametric regression procedure to estimate the infinite-dimensional conditional likelihood bounds — this is captured by the oracle’s integrated MSE. The second is estimation error from first-stage nuisance function estimation. Our next result shows that the latter only affects our proposed estimator through a smoothed, doubly-robust remainder term.
@@
Let $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ denote the second-stage pseudo-outcome regression estimator. Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption A.1), and $\mathbb{P}(\epsilon\leq\hat{\pi}_{1}(X_{i})\leq 1-\epsilon)=1$ for some $\epsilon>0$. Define $\tilde{R}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, and $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$. Then,
@@
Proposition 3.1 delivers two insights. First, our analysis is modular: researchers can pair choices of nuisance function estimators in the first step with any choice of nonparametric regression method (satisfying Assumption A.1) in the second step, and the integrated MSE is characterized by the oracle rate plus a doubly-robust remainder term. Second, Proposition 3.1 shows that the cost of estimating the nuisance functions in the first step only enters through the product of nuisance function errors.
To prove this result, we establish an oracle inequality on the $L_{2}(\mathbb{P})$-error of pseudo-outcome regression procedures (Lemma A.1), extending Kennedy(22)-towardsoptimal’s pointwise analysis. This result may be of independent interest. The $L_{2}(\mathbb{P})$-stability condition (Assumption A.1) on $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ is quite mild in practice. It is satisfied by a variety of generic linear smoothers such as linear regression, series regression, nearest neighbor matching, random forest models, and several others (Proposition A.1). Once again, in this sense, Proposition 3.1 is agnostic to both the researcher’s choice of $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$} and nuisance function estimators.
[…]
@@ 3.2 Regret Bound for Plug-In Decision Rules
Researchers often apply a threshold rule to an estimated predictive algorithm to make decisions. For example, KLLLM(18); Rambachan(21); AngelovaEtAl(22) compare outcomes under alternative decisions rules that threshold a predictive algorithm in pretrial release. We show that plug-in decision rules based on our estimators of the conditional likelihood bounds achieve low regret, ensuring that the statistical uncertainty from estimation does not lead to substantially worse decisions than if the true conditional probabilities were known.
Since the expected payoff $U(d)$ of any decision rule $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ is partially identified (Lemma 2.2), we evaluate decision rules by comparing their worst-case expected welfare (e.g., Manski(07)). An immediate consequence of Lemma 2.2 is that the optimal max-min decision rule $d^{*}(\cdot)\in\arg\max_{d(\cdot)\colon\mathcal{X}\rightarrow[0,1]}\underline{U}(d)$ is a threshold rule:
$d^{*}(X_{i})=1\{\widetilde{\mu}^{*}(X_{i})\leq u_{1,0,i}+u_{0,0,i}\},$
[…]
@@
$R(\widehat{d})^{2}\leq 2\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|+2\|\widehat{\underline{\mu}}_{oracle}(\cdot)-\underline{\mu}^{*}(\cdot)\|+4\epsilon^{-1}\sqrt{2}\|\tilde{R}(x)\|+o_{\mathbb{P}}(R_{oracle}).$
The proof proceeds in two steps. First, we show that the regret of the plug-in decision rule is bounded by the integrated MSE of our estimators for the conditional probability bounds — this follows because the optimal decision rule is a threshold rule, and threshold rules are sensitive to estimation errors measured in $L_{2}(\mathbb{P})$ norm. Second, we apply Proposition 3.1 to decompose this integrated MSE into the oracle rate plus the doubly-robust remainder term. Consequently, worst-case regret converges to zero whenever the oracle’s integrated MSE converges to zero. Plug-in decision rules based on our nonparametric estimators are therefore robust: they achieve nearly optimal welfare despite estimation uncertainty.
@@ 4 Robust Evaluation of Predictive Algorithms
In this section, we construct estimators for the bounds on the performance of a given predictive algorithm $s(\cdot)$ under Assumption 2.1. This setting arises empirically in two scenarios: (i) when researchers evaluate an externally provided algorithm (e.g., a proprietary credit score or a commercial risk assessment tool), and (ii) when researchers conduct hold-out evaluation of an algorithm trained on separate data. In both cases, the algorithm $s(\cdot)$ is taken as fixed, and the goal is to assess its performance under varying assumptions about unobserved confounding. Our estimators are consistent for the bounds on the identified set as sample size grows large, and we characterize their convergence rates in terms of errors in the first-step estimation of nuisance parameters.
4.1 Bounds on Overall Performance
We first construct estimators for the bounds on overall performance $\mbox{perf}(s;\beta)$. To simplify exposition, we develop estimators for observed outcome bounds (Section 2.3.1). We extend to proxy outcome bounds and instrumental variable bounds in Online Appendices D.1-D.2 respectively. Under observed outcome bounds, Lemma 2.1 implies that the upper bound can be written as
The lower bound $\underline{\mbox{perf}}(s;\beta)$ has an analogous expression. Both are linear functionals of known functions of the data and identified nuisance parameters $\eta=(\mu_{1},\pi_{0})$. This linearity enables us to construct debiased estimators using standard arguments based on influence functions and cross-fitting (e.g., ChernozhukovEtAl(18)-DML; ChernozhukovEtAl(22)-LocallRobust; Kennedy(22)-IFreview).
Estimation Procedure for Observed Outcome Bounds:
For completeness, we sketch the construction of our estimators based on $K$-fold cross-fitting. We randomly split the data into $K$ disjoint folds. For each fold $k$, we estimate the nuisance functions $\widehat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
for each observation $i$ in the $k$-th fold, where $\phi_{\mu,i}(\eta)$ and $\phi_{\pi\mu,i}(\eta)$ are defined as before. We then average over all observations $\widehat{\overline{\mbox{perf}}}(s;\beta):=\mathbb{E}_{n}\left[\overline{\mbox{perf}}_{i}(\hat{\eta}_{-K_{i}})\right],$ where $K_{i}$ is the fold containing observation $i$. Our estimator for the lower bound $\widehat{\underline{\mbox{perf}}}(s;\beta)$ is defined analogously.
Our next result characterizes the convergence rate and asymptotic distribution of our estimators for the bounds on overall performance under observed outcome bounds.
Proposition 4.1.
Define $R_{1,n}^{k}:=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$ for each fold $k$. Assume (i) there exists $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$, (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$ for each fold $k$, and (iv) $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$ for each fold $k$. Then,
$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,\Sigma\right)$
The estimators converge at rate $O_{\mathbb{P}}(1/\sqrt{n})$ plus a term that captures the nuisance estimation error. Nuisance errors enter multiplicatively — through the product of propensity score and outcome regression errors — rather than additively. The rate condition required for the estimators to be jointly asymptotically normal is satisfied under the standard condition that all nuisance function estimators converge at a rate faster than $O_{\mathbb{P}}(n^{-1/4})$. This is the familiar condition from debiased machine learning, satisfied by a wide range of nonparametric regression and modern machine learning methods.
Online Appendix C.2 constructs a consistent estimator of the asymptotic covariance matrix in Proposition 4.1. This implies that confidence intervals for the identified set $\mathcal{H}(\mbox{perf}(s;\beta))$ can be constructed using standard methods (e.g., ImbensManski(04); Stoye(09)--PICI).
@@ 4.2 Bounds on Positive Class Performance
We next estimate bounds on positive class performance $\mbox{perf}_{+}(s;\beta)$. Unlike overall performance, this is a ratio of expectations, and the bounds are given by the optimal values of linear fractional programs as shown in Lemma 2.1. This creates additional challenges in estimation. In this section, we will develop estimators for these bounds and establish their convergence properties.
4.2.1 Estimation via Linear Fractional Programming:
[…]
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\tilde{\delta}\in\widehat{\Delta}_{n}}\frac{\mathbb{E}_{n}\left[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}\right]}{\mathbb{E}_{n}\left[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}\right]},$ (7)
[…]
We provide an error decomposition for the positive class performance bounds. This result characterizes how errors in estimating the nuisance parameters $(\mu_{1}(\cdot),\pi_{1}(\cdot))$ and the bounding functions $(\underline{\delta}(\cdot),\overline{\delta}(\cdot))$ propagate to the final estimator.